OpenStudy (anonymous):

Four-fifths of a jar is filled with juice. Three children split the juice equally. How much juice will each child get?

OpenStudy (lgbasallote):

you simply divide 4/5 by 3 \[\frac 45 \div 3 \implies \frac 45 \times \frac 13 \implies \frac{4}{5 \times 3}\] do you get it?